OpenStudy (anonymous):

simpliy x2-10x+9/ x2+2x-3

OpenStudy (lgbasallote):

factor the numerator and denominator..do you know how to?

OpenStudy (anonymous):

=(x-1)(x-9) / (x-1)(x+3)=(x-9)/(x+3) Im not sure if thats right but @lgbasallote will explain you more about it..

OpenStudy (lgbasallote):

yes that's right :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No Problem.

OpenStudy (anonymous):

@lgbasallote wait is it -2ac6?

OpenStudy (lgbasallote):

where did c and 6 come from again?

OpenStudy (lgbasallote):

and a?

OpenStudy (anonymous):

@lgbasallote i didnt mean to say that lol sorry.

OpenStudy (lgbasallote):

so do you know the answer to your problem now?